Journalist Report – February 14th

Author: Maria Grulich

Title: Love, Bacon and Cookies – It’s Valentine’s Day

Today is a very special day for the crew! It’s the start of the leadership rotation and also it is Valentine’s Day!

The crew has prepared some cookies and our new commander and XO, Dave and Nathan prepared everyone a breakfast of eggs and bacon as a surprise for the crew. After watching The Martian last night, the crew is thinking on how Matt Damon solved the persistent problems with living on Mars. Now the crew is thinking how to germinate potatoes here!

Currently our most valuable resource is data. We solved the water problem by changing the way we clean dishes, so we are on one refill a day. Regarding the data, we still have to investigate the anomaly.

Today only one EVA was planned with the objective of flying the drone for the first time. The preparation went smoothly since we have fixed and inventoried the EVA suits. The crew composed of Natalia (GHO), Daniel (HSO), Veronica (ENG) and Maria (Media) stepped out at 1030 and the weather was cloudy, but stable with a slight wind. Directly after leaving the Hab, the drone was tested to see how it performs and it worked fine. After a couple of tests flights, it was decided to keep moving to a higher altitude to take pictures with the drone from there. After 10 more minutes in the weather conditions changed and it was decided to abort the EVA. It seems like the North Ridge is our curse. 3 EVAs planned there and three cancelations. The crew is still very determined to explore the North Ridge during this rotation.

After the short EVA, the crew found themselves with some free time which gave us the opportunity to exchange our Secret Cupid gifts. Every crew member brought a small gift for another crew member and it was exciting to see who picked whom and it was accompanied with heart shaped Martian cookies.

Today also gave time to start for the video project as well as the detailed categorization and analysis of the rock samples. On one of the rocks that was brought back from Candor Chasma, a skeleton was found. It is probably a skeleton from a scorpion which is a really interesting finding.

Tomorrow, there will be time for more indoor activities and outreach activities.

Ad Astra!

GreenHab Report – February 14th

GreenHab Officer: Natalia Larrea

Environmental control: Heating and Cooling w/ ambient air (8 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 17.5 °C

High: 26.5°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 12 gallons

Water in Blue Tank: ~ 252.5 gallons

Time(s) of watering for crops:

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)
[AT TIME OF WATERING]

1) 09.00 AM 24.5° C / 28 %
2) 12.00 PM 25.2° C / 32 %
3) 14:40 PM 24.5° C / 38 %
4) 18:00 PM 20.6° C / 48 %

Changes to crops: New seedlings continue sprouting around the large tomato plant in the grow bed. Some plants exhibit potential nutrient deficit (see narrative below). The A. Thaliana started sprouting on the germination plate.

Narrative: Further seeds of A. Thaliana were washed and placed in germination plates using damp paper towel germinations last night (13th February). The plates were placed inside the incubator to maintain optimal environmental conditions. The A. Thaliana that was plated on Monday started sprouting as well and will be transferred to collected regolith samples tomorrow. We continued regolith characterization by investigating the pH of the sample. The samples were treated with a CaCl2-2H2O solution and de-ionized water results will be available in tomorrow’s report.

The GreenHab temperature was regulated throughout the day with the use of the heater at specific intervals of time. We conducted inventory in the GreenHab to identify which plants exhibit yellow pigmentation (i.e., possible indicator of nitrogen/nutrient deficiency in those pots). The following plants were identified as having a nutrient deficit (see attached pictures):

· Savory mix (some leaves started to turn yellow)

· Lettuce encore mix (some leaves exhibit slight yellow color)

· Lettuce leaf (some leaves started to turn yellow)

· Parsley (the plant is largely healthy, few leaves at the bottom started turning yellow).

· Fennel (the plant is largely healthy, few leaves at the bottom started turning yellow).

· Sugar snap peas (most plants turning yellow. Some drying)

· Cucumber (plants seem to be dead, as previously reported)

We prepared the plants nutrient employing the “Miracle-Gro” solution available in the GreenHab. We sprayed the solution sparingly at specific spots at the base of the affected plants. We will continue monitoring the situation to observe how plants react. It is anticipated we will repeat this process in the next two to three days.

Harvest: 2g parsley, 1g lemon basil, 1g fennel greens.

Support/supplies needed: None.

Operations Report – February 14th

SOL: 4

Name of person filing report: Verónica Triviño

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator (hours run): Turned on at 1730 (14th Feb), turned off at 1500 (14th Feb)

Solar— SOC % – Turned off Gen (14th Feb) 83%, Turned on Gen (14th Feb) 67%

Diesel Reading – 97%

Propane Reading – ~ 50%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – around 68%, ~ 372 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 252.5 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01412244 units

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Not Used

Hours: 50.1 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not Used

Hours: 74.3 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 73.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: N/A

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Yes, going to town

Crew Car used and why, where? Not used

General notes and comments: Nothing to report

Summary of internet: Around noon, D. Masaitis (CO) checked the internet data status and saw a remaining data allocation of 392 MB. 20 minutes later, the data allocation dropped to 0 MB. Upon inspection of the crew’s devices, no offending device was found. A survey of the devices connected to the internet revealed only the router and D. Masaitis’ laptop, which was being used to correspond with Mission Control. The cause of the anomaly is still unknown and we will continue to monitor the Hab network.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Temperature in the upper deck of the Hab rose during the morning, at some points getting close to reaching 27 deg C, even though the heater in the lower deck was set by the engineer to slightly below the "ideal" middle setting of "3" (which usually gives the Hab a temperature of 21 – 23 deg C). The assistant director recommended turning the heater off and on again. After turning the heater off, the engineer and commander struggled to turn it back on again. As of writing this report, it is currently still off, however the temperature in the upper deck is a comfortable 23.8 deg C.

EVA Report – February 14th

EVA #07

Author: N. Hadland, Executive Officer

Purpose of EVA: Terrain familiarization and sample collection

Start time: 1030

End time: 1115

Narrative: The crew left the Hab at approximately 10:35 and headed North towards North Ridge. In a clear space about 200m North of the robotic observatory they tested their drone “Parrot Mambo No. 5” for the first time. It succeeded in taking off with its ribbon attachments below but was noticeably perturbed in its flight by the winds so it landed and the ribbons were removed. The ribbon’s purpose is part of a proof of concept experiment to see whether a hovering drone with colorful ribbon below can be used as a beacon or lighthouse for nearby lost EVA crews. The drone took off a second time without this attachment and flew to an appreciable height of approximately 10m before the wind began blowing the drone too strongly and it was landed to prevent damage to the equipment. It was then put on charge on a portable power bank and the crew set off North again along Path 1103. Approximately 300m from the base of North Ridge in a hilly area the drone was flown again to take photographs of the crew below with the vertical camera. 4 photos were taken, some of which successfully show the crew below with good resolution. Before the crew began their approach to the North Ridge ascent on Path 1103 at approximately 1055, COMMS Officer H. Blackburn called in to notify the crew that the incidental weather they had been monitoring had reached the Hab and precipitation would likely reach the EVA team in under 10 minutes. For safety, the EVA was stopped and the crew returned to the Hab before the rain made the ground unsuitable for exploration. Effective monitoring of the environmental conditions at the Hab and clear communication with the EVA team ensured this was a productive and safe short EVA. The EVA team returned to the airlock at 1110 without incident.

Destination: Pooh’s Corner, North Ridge

Coordinates: 0518800/4250900 (Pooh’s Corner), 0518600/4251500 (North Ridge)

Participants: Natalia Larrea (GHO), Daniel Robson (HSO), Maria Grulich (MEDIA) Veronica Trivino (ENG)

Roads and routes per MDRS Map: North along Cow Dung Road (0110), through Pooh’s Corner, and then northwest to North Ridge

Mode of Travel: Walking

Sol Summary – February 14th

Sol: 4

Summary Title: Nothing Says Love Like Bacon!

Author’s name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: Crew 205 awoke to overcast skies and the smell of bacon and eggs, as the Commander and Executive Officer cooked a breakfast of bacon, eggs, and toast prepared from homemade bread. Unfortunately, the only sunny things today were the egg yolks. The EVA team for EVA #07 left on schedule, but was recalled 45 minutes later as weather began to turn inclement. The EVA crew was in the airlock before any unsafe ground conditions occured, but they only completed the UAV test flight, and did not make it to the North Ridge. The rest of the crew helped the EVA team clean and store their EVA suits, before everyone headed upstairs for lunch. While having lunch, the crew engaged in a "Secret Valentine" gift exchange and then used the afternoon to catch up on scientific analysis of samples collected over the week’s EVAs so far. Towards the evening, the crew gathered in the upstairs common area of the Hab to prepare reports for the communications window, share stories, and discuss what would be prepared for dinner. After the communications window, the crew will eat chicken and rice casserole, discuss tomorrow’s priorities of work, and then play some card games before bed.

Look Ahead Plan: Due to today’s ongoing inclement weather, the ground tomorrow will be too saturated to make EVAs practical. Instead, the crew will remain at the Hab to focus on refining research, improving ongoing investigations, and drafting procedural documents to improve the workflow of future crews. The inability to conduct EVAs can be frustrating, but having multiple crew members with culinary skills makes being stuck at the Hab much more palatible. Tomorrow will also see another rotation of Commander and Executive Officer, so a new day will bring new leadership.

Anomalies in work: EVA# 07 terminated early due to weather

Weather: Cloudy all day long, rain started after 1100 hrs, with consistent rain all afternoon and into the evening

Crew Physical Status: Nominal

EVA:

EVA#07: 0518800/4250900 (Pooh’s Corner), 0518600/4251500 (North Ridge, not reached)

Reports to be filed: Sol Summary, EVA report, Operations Report, Greenhab Report, Journalist Report

Support Requested: None

Journalist Report – February 13th

Crew 205 Crew Journalist Report 13-FEB-2019 Sol 2

Author: Maria Grulich

Title: Water, Water, Everywhere

After our Spanish night yesterday, the girls decided to open the first water-efficient hair wash salon on Mars. It felt like on Earth we just had to skimp on the extra products 🙂

After this, the crew woke up refreshed and ready to tackle another day on Mars, including another 4 hour EVA. Dave, Hannah, Veronica, and Nathan left at 10:25 to go towards Candor Chasma. Exploring the cannon, they found beautiful rocks which seemed to be perfectly smooth, a sign of past liquid water. Dave and Nathan seemed to be bighorn sheep in their former life climbing up the Martian mountains.

The team achieved all its objectives and even had time to take some pictures.

The Hab team checked all EVA suits and replaced several batteries to make sure that all EVA suits are working properly before going out.

At 15:00 the second EVA had the plan to go up the mountain and take some samples and fly the drone to take pictures of the terrain. Unfortunately, the wind picked up and we decided not to take out the drone. Still, we planned to collect samples of stratified layers.

Once we stepped out and climbed the first hills the weather changed, and the team could only collect one sample out of the planed 4 of the stratified layers.

The team returned after 50 min. The fog inside the helmet made it very hard to climb down the mountain. Still, the day will be used to do an analysis of the samples, to prepare astronomy observation, maintain the Hab, water the GreenHab, and start the first videos about life on Mars. We also met yesterday’s goal of using only one loft tank of water.

One lesson learned is that data and water are currently our most precious resources and more measures to protect these resources have to be taken.

After the work is done and all reports are submitted the crew 205 will have a movie night.

Ad Astra!

Greenhab Report – February 13th

Crew 205 Green Hab Report – 13-FEB-2019

GreenHab Officer: Nathan Hadland

Environmental control: Heating and Cooling w/ ambient air (0 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 17.4°C

High: 27.3°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 12 gallons

Water in Blue Tank: ~ 264.5 gallons

Time(s) of watering for crops:

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)
[AT TIME OF WATERING]

1) 09.00 AM 21.4° C / 28 %
2) 12.00 PM 22.2° C / 24 %
3) 15:30 PM 21.9° C / 27 %
4) 18:15 PM 25.5° C / 33 %

Changes to crops: New seedlings sprouting around the large tomato plant in the grow bed.

Narrative: We continued regolith sample collection in Candor Chasma and will continue characterization throughout the evening using the dissection and optical microscopes. We are also performing spectral analysis using a spectrometer for preliminary elemental characterization.

The A. Thaliana has not germinated yet on the agar germination plates, so we are considering using a damp paper towel germination as a backup. The GreenHab dropped in temperature today due to persistent cloud cover. However, the heater mitigated the temperature fluctuations.

The red leaf lettuce us not showing any pigmentation, which is a possible indicator of plant stress. Additionally, the peas still have yellow pigmentation, which leads us to believe there are nutrient and nitrogen deficiency in those pots. We will continue to monitor these experiments.

We are still waiting for the arrival of seeds.

Harvest: None.

Support/supplies needed: None

Science Report – February 13th

Science Report 13 February 2019

Crew 205 – International Emerging Space Leaders

Submitted and prepared by GreenHab Officer Nathan Hadland

1.) ISRU Study: Geological characterization of regolith samples began with observations under the optical and dissection microscopes. We have also been performing spectral analysis using a spectrometer to get the initial elemental characterization. Analysis of pH and redox potential of samples to determine the suitability of the substrate to host plant growth will be investigated in the coming days. Additionally, the Arabidopsis thaliana seeds were prepped using a washing protocol and plated on an agar germination plate. We are waiting on germination to occur. Dead plant material from the GreenHab was taken and dried using an oven in the Science Dome and subsequently weighed for the cellulosic ethanol ISRU study.

2.) GreenHab: Dead and live samples of cucumbers were taken from the GreenHab and wet mounts were prepared and viewed under the optical microscope. We found evidence of degraded cell structure, likely from the fungi growing on the dead plants. The live tissue comparatively had all of the expected components of plant cells including stoma and clear cell walls. The carrots that were not harvested were moved from around the grow bed and consolidated into one section of the planter.

3.) Astronomy: The variable star SY MON was observed on 11 Feb 2019. The observation session was one 60 s exposure using the V filter. On 12 Feb 2019, the magnitude value was measured for SY MON using AstroImage J software. The magnitude value was found to be 14.147. The value will be submitted to the AAVSO. Another observation session was submitted on 12 Feb 2019. The session was for the ASASSN-V J081823.00-111138.9 variable star with the aim of drawing a light curve. The observation session was 9 exposures in the V filter and 9 exposures in the B filter. The V filter exposures were 60 seconds, and the B filters exposures were 120 seconds. However, due to the bad weather, no images were captured.

4.) EVA Project: The primary research technique was changed after testing analysis methods from the mission plan during EVA’s of the first 2 days. A new research focus has been adopted and is currently under development. In practice, a “score” will be assigned to each EVA to determine its effectiveness by evaluating various parameters. More data will be available in upcoming reports.

5.) Effective Leadership: Role rotations will begin tomorrow and the first survey conducted on the current commander, N. Larrea will be collected tonight.

6.) Developing Guidelines (SOPs): Water conservation procedures, improved communications procedures, and EVA preparation procedures are being developed. An inventory of the Science Dome, GreenHab, and the RAM are also being compiled as well as an improved list of food provided to crews.

Astronomy Report – February 13th

Crew 205 Astronomy Report 13 Feb 2019

Crew Astronomer: Ghanim Alotaibi

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested: MDRS-14

Objects Viewed: Out of the 18 images requested last night, I found only 12 images were captured. However, all the captured images were found to be bad for photometry. The seeing condition was not good yesterday for astrophotography and it was decided to resubmit the observation session.

Problems Encountered: Bad weather conditions.